# Exact computation of Transfer Entropy with Path Weight Sampling

### v3

@@ Abstract
The ability to quantify the directional flow of information is vital to understanding natural systems and designing engineered information-processing systems. A widely used measure to quantify this information flow is the transfer entropy. However, until now, this quantity could only be obtained in dynamical models using approximations that are typically uncontrolled. Here we introduce a computational algorithm called Transfer Entropy-Path Weight Sampling (TE-PWS), which makes it possible, for the first time, to quantify the transfer entropy and its variants exactly for any stochastic model, including those with multiple hidden variables, nonlinearity, transient conditions, and feedback. By leveraging techniques from polymer and path sampling, TE-PWS efficiently computes the transfer entropy as a Monte-Carlo average over signal trajectory space. We apply TE-PWS to linear and nonlinear systems to reveal how transfer entropy can overcome naive applications of the data processing inequality in the presence of feedback.
@@
However, there are currently no exact methods to compute the transfer entropy in a general many-variable dynamical model. The transfer entropy depends on probability distributions of signal trajectories. Estimating these distributions by binning the experimental or simulated trajectories in histograms is not feasible, as the dimensionality of the trajectory space scales exponentially with signal duration 9, 10. As a result, different approximations are currently being used: either the full history dependence of the transfer entropy is truncated11, or an arbitrary distance metric in trajectory space is chosen for clustering the trajectories12, or a linear or low-order moment-closure approximation is employed13, 14, 15. These approximations can result in uncontrolled errors in complex, nonlinear, many-variable systems16, 17.
Here we fill this gap by introducing TE-PWS, a numerical algorithm to estimate transfer entropies exactly for any stochastic model, including diffusive and jump processes. The estimate is exact, i.e., it is an unbiased statistical estimate of the transfer entropy. TE-PWS can therefore provide ground truth results for any given model. TE-PWS builds on the recently developed PWS algorithm for computing the mutual information between trajectories18. TE-PWS exploits the idea that path likelihoods can be obtained analytically from the Langevin or master equation, from which the transfer entropy is then computed via Monte-Carlo averaging in trajectory space. Additionally, long trajectories are sampled with an importance sampling scheme19, solving the problem of exponential scaling of the computational cost with trajectory duration. We have applied TE-PWS to compute the transfer entropy in a three-variable motif in the presence of feedback, for both linear and nonlinear systems. We show that TE-PWS reproduces analytical results when available and produces novel insights on how information feedback can amplify information transfer. Specifically, the transfer entropy from an input to an output node can overcome a naive application of the data processing inequality even when the mutual information obeys one.
Diffusive process. We introduce transfer entropy and TE-PWS first in the context of diffusive processes. Consider a $d$-dimensional diffusive process $\mathbf{X}(t)$ modelled as a function of time $t$ by a Langevin equation
@@
where the index $k$ goes over individual timesteps, $X_{j}(k)$ denotes $X_{j}$ after $k$ timesteps, $X_{j,[0,k]}$ denotes the trajectory over the first $k$ timesteps, and $I(A;B)$ denotes the mutual information between two random variables $A$ and $B$. $\mathcal{T}_{X_{i}\to X_{j}}$ measures the information transferred from the past trajectory of $X_{i}$ to the new updates of $X_{j}$ at every timestep, given the past trajectory of $X_{j}$ is already known. In case the dynamics relaxes into a steady-state, we will also talk about the transfer entropy rate, $\dot{\mathcal{T}}_{X_{i}\to X_{j}}=\lim_{N\to\infty}\mathcal{T}_{X_{i}\to X_{j}}/(N\delta t)$.
@@
$\mathcal{T}_{X_{i}\to X_{j}}=$ $\sum_{k}\Delta\mathcal{T}^{i\to j}_{[k,k+1]}=\sum_{k}H\left(X_{j}(k+1)|X_{j,[0,k]}\right)$
$\qquad-H\left(X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]}\right)$ (3)
@@
where $\Delta\mathcal{T}^{i\to j}_{[k,k+1]}$ denotes the increment at every timestep, $H(A)$ denotes the Shannon entropy associated with the probability distribution $P(A)$ of $A$, and the angular brackets denote an average over the joint probability $P\left(X_{i,[0,N]},X_{j,[0,N]}\right)$. Eq. 3 shows that transfer entropy quantifies the additional information in $X_{j}(k+1)$ that arrives from $X_{i,[0,k]}$ beyond that which is already present in the past trajectory $X_{j,[0,k]}$. This occurs either through direct causal action, or through a third variable $X_{l}$, schematically demonstrated in Fig. 1a. If $X_{i}$ does not affect the dynamics of $X_{j}$, this additional information would be zero. In general, every $\Delta\mathcal{T}^{i\to j}_{[k,k+1]}$ is a mutual information and hence nonnegative.
@@
where the index $\nu$ sums over $M_{1}$ pairs of trajectories of $X_{i}$ and $X_{j}$ sampled from the joint probability distribution $P\left(X_{i,[0,N]}^{(\nu)},X_{j,[0,N]}^{(\nu)}\right)$. For each pair of trajectories, the probabilties in the numerator and the denominator of Eq. 5 are not analytically available, but what is indeed available is the full joint probability $P(\mathbf{X}_{[0,N]}^{(\nu)})$ as the exponential of the Onsager-Machlup action20, 21. We thus need to marginalize over all degrees of freedom other than $X_{i}$ and $X_{j}$, denoted henceforth collectively as $X_{l}$. We illustrate this procedure first for the denominator in Eq. 5. It is obtained as $P(X^{(\nu)}_{j,[0,k+1]})/P(X^{(\nu)}_{j,[0,k]})$, where $P(X^{(\nu)}_{j,[0,k]})$ is obtained via marginalization,
@@
where the index $\mu$ sums over $M_{2}$ trajectories sampled from $P_{0}$. What is the best choice for $P_{0}$? The ideal choice would be the conditional distribution $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]}^{(\nu)})$, as it makes the summand in Eq. 7 equal for all $\mu$, such that the variance of the estimate of $P(X_{j,[0,k]}^{(\nu)})$ is zero. However, this conditional distribution is not known a priori. We therefore generate $X_{i}$ and $X_{l}$ trajectories in the frozen field of $X_{j,[0,N]}^{(\nu)}$ (Fig. 1b) resulting in a distribution $P_{0}$ that is known analytically and is expected to be close to this conditional distribution. To exactly compensate for the remaining deviations of $P_{0}$ from the desired conditional distribution, we employ, in the spirit of Rosenbluth-Rosenbluth(RR)-PWS18, a reweighing of the $X_{i}^{(\mu)}$ and $X_{l}^{(\mu)}$ trajectories on-the-fly with weights proportional to the ratio of the two distributions. We resample the trajectories with these weights after every $\delta t$ time, meaning that we sample $M_{2}$ trajectories with replacement from the weighted trajectory ensemble (Fig. 1c). This procedure exactly generates $X_{i}^{(\mu)}$ and $X_{l}^{(\mu)}$ trajectories according to the desired conditional distribution $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]}^{(\nu)})$ (see End Matter (EM) and Supplemental Material (SM)).
@@
To summarize, for each of the $M_{1}$ pairs of $X_{i,[0,N]}$ and $X_{j,[0,N]}$ trajectories, we simulate a joint ensemble of $M_{2}$ new $X_{i}$ and $X_{l}$ trajectories to estimate the denominator in the logarithm in Eq. 5, and a separate ensemble of $M_{2}$ new $X_{l}$ trajectories to estimate the numerator (Fig. 1). The trajectories in each ensemble are resampled on-the-fly with the RR scheme, giving both the numerator and denominator in Eq. 5 as Monte-Carlo averages. The computational cost thus scales as $2M_{1}M_{2}$. The transfer entropy estimate is unbiased 18 and the statistical accuracy can be arbitrarily improved by increasing $M_{1}$ and $M_{2}$. A pseudocode for the algorithm is available in the SM. Aside from Schreiber’s transfer entropy1, other trajectory-based metrics of directional information transfer such as directed information22, conditional transfer entropy23 and filtered transfer entropy24, also can be derived from conditional distributions of trajectories. Hence TE-PWS can be used to compute all such metrics at similar cost, as shown in the SM.
@@
We demonstrate the validity of the method by analyzing two simple examples of an Ornstein-Uhlenbeck (OU) process for which transfer entropy rates are analytically available. Eq. 1 is an OU process when it is linear with $\mathbf{F}=-\mathbf{a}\mathbf{X}$ where $\mathbf{a}$ is a spring constant matrix. The first example is a minimal three-dimensional model for describing the stochastic dynamics of gene expression and the growth rate of bacteria25. The model is non-bipartite, meaning that the diffusion constant matrix is non-diagonal and the mutual information rates are not finite. Yet transfer entropy rates are finite and semi-analytically available24. The other example is an OU process with two variables, for which transfer entropy rates are analytically available (see EM). As shown in Fig. 2a and b, TE-PWS results converge to the known transfer entropy rates within a few multiples of the relaxation timescales in the respective systems. Details about the models are provided in the EM.
@@
Figure 2: Convergence of transfer entropy rate estimates. (a) and (b) respectively show forward and reverse transfer entropy rates in the stochastic gene expression model in units of an inverse relaxation timescale, $\mu_{0}^{-1}$, as a function of the trajectory duration $T$. Symbols are from TE-PWS and dashed lines are numerically exact values. (c) and (d) show transfer entropy rates in the two-variable OU process. Symbols are from TE-PWS and dashed lines are analytical results.

### v4

@@ Abstract
The ability to quantify the directional flow of information is vital to understanding natural systems and designing engineered information-processing systems. A widely used measure to quantify this information flow is the transfer entropy. However, until now, this quantity could only be obtained in dynamical models using approximations that are typically uncontrolled. Here we introduce a computational algorithm called Transfer Entropy-Path Weight Sampling (TE-PWS), which makes it possible, for the first time, to quantify the transfer entropy and its variants exactly for any stochastic model, including those with multiple hidden variables, nonlinearity, transient conditions, and feedback. By leveraging techniques from polymer and path sampling, TE-PWS efficiently computes the transfer entropy as a Monte-Carlo average over signal trajectory space. We use our exact technique to demonstrate that commonly used approximate methods to compute transfer entropies incur large systematic errors and high computational costs. As an application, we use TE-PWS in linear and nonlinear systems to reveal how transfer entropy can overcome naive applications of the data processing inequality in the presence of feedback.
@@
However, there are currently no exact methods to compute the transfer entropy in a general many-variable dynamical model. The transfer entropy depends on probability distributions of signal trajectories. Estimating these distributions by binning the experimental or simulated trajectories in histograms is not feasible, as the dimensionality of the trajectory space scales exponentially with signal duration9, 10. As a result, different approximations are currently being used: either the full history dependence of the transfer entropy is truncated11, or an arbitrary distance metric in trajectory space is chosen for clustering the trajectories12, or a linear or low-order moment-closure approximation is employed13, 14, 15. These approximations can result in uncontrolled errors in complex, nonlinear, many-variable systems16, 17. In the absence of an exact method, the magnitude of these errors remains unclear.
Here we fill this gap by introducing TE-PWS, a numerical algorithm to estimate transfer entropies exactly for any stochastic model, including diffusive and jump processes. The estimate is exact, i.e., it is an unbiased statistical estimate of the transfer entropy. TE-PWS can therefore provide ground truth results for any given model. TE-PWS builds on the recently developed PWS algorithm for computing the mutual information between trajectories18. TE-PWS exploits the idea that path likelihoods can be obtained analytically from the Langevin or master equation, from which the transfer entropy is then computed via Monte-Carlo averaging in trajectory space. Additionally, long trajectories are sampled with an importance sampling scheme19, solving the problem of exponential scaling of the computational cost with trajectory duration. We first show that TE-PWS reproduces analytical results when available. We then apply TE-PWS to compute the transfer entropy in a three-variable motif in the presence of feedback, for both linear and nonlinear systems, yielding novel insights on how information feedback can amplify information transfer. Specifically, the transfer entropy from an input to an output node can overcome a naive application of the data processing inequality even when the mutual information obeys one. Finally, we use TE-PWS to show that the most widely used approximate methods for the computation of transfer entropy— the Gaussian framework, KSG and single-step truncation— all exhibit significant systematic errors. The comparison also reveals that TE-PWS is computationally efficient.
Transfer entropy. The TE-PWS algorithm has the same key steps for all stochastic models. We therefore describe TE-PWS for diffusive processes in the main text and for jump processes in the End Matter (EM). Consider a $d$-dimensional diffusive process $\mathbf{X}(t)$ modelled as a function of time $t$ by a Langevin equation
@@
$\equiv\sum_{k=0}^{N-1}\mathcal{T}_{X_{i}\to X_{j}}^{[k]}$ (3)
where the index $k$ goes over individual timesteps, $X_{j}(k)$ denotes $X_{j}$ after $k$ timesteps, $X_{j,[0,k]}$ denotes the trajectory over the first $k$ timesteps, $\mathcal{T}_{X_{i}\to X_{j}}^{[k]}$ denotes the $k$-th term in the sum, and $I(A;B)$ denotes the mutual information between two random variables $A$ and $B$. $\mathcal{T}_{X_{i}\to X_{j}}$ measures the information transferred from the past trajectory of $X_{i}$ to the new updates of $X_{j}$ at every timestep, given the past trajectory of $X_{j}$ is already known. In case the dynamics is in steady-state, we will also talk about the transfer entropy rate, $\dot{\mathcal{T}}_{X_{i}\to X_{j}}=\lim_{N\to\infty}\mathcal{T}_{X_{i}\to X_{j}}/(N\delta t)$.
@@
$\mathcal{T}_{X_{i}\to X_{j}}$ $=\sum_{k=0}^{N-1}\mathcal{T}_{X_{i}\to X_{j}}^{[k]}=\sum_{k}H\left(X_{j}(k+1)|X_{j,[0,k]}\right)$
$\qquad\qquad-H\left(X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]}\right)$ (4)
@@
where $H(A)$ denotes the Shannon entropy associated with the probability distribution $P(A)$ of $A$, and the angular brackets denote an average over the joint probability $P\left(X_{i,[0,N]},X_{j,[0,N]}\right)$. Eq. 4 shows that transfer entropy quantifies the additional information in $X_{j}(k+1)$ that arrives from $X_{i,[0,k]}$ beyond that which is already present in the past trajectory $X_{j,[0,k]}$. This occurs either through direct causal action, or through a third variable $X_{l}$, schematically demonstrated in Fig. 1a. If $X_{i}$ does not affect the dynamics of $X_{j}$, this additional information would be zero. In general, every $\mathcal{T}_{X_{i}\to X_{j}}^{[k]}$ is a mutual information and hence nonnegative.
@@
where the index $\nu$ sums over $M_{1}$ pairs of trajectories of $X_{i}$ and $X_{j}$ sampled from the joint probability distribution $P\left(X_{i,[0,N]},X_{j,[0,N]}\right)$. The superscript $(\nu)$ in $X_{i}^{(\nu)}$ refers to the $\nu$-th statistical realization of $X_{i}$. For each pair of trajectories, the probabilties in the numerator and the denominator of Eq. 6 are not analytically available, but what is indeed available is the full joint probability $P(\mathbf{X}_{[0,N]}^{(\nu)})$ as the exponential of the Onsager-Machlup action20, 21. We thus need to marginalize over all degrees of freedom other than $X_{i}$ and $X_{j}$, denoted henceforth collectively as $X_{l}$. We illustrate this procedure first for the denominator in Eq. 6. It is obtained as $P(X^{(\nu)}_{j,[0,k+1]})/P(X^{(\nu)}_{j,[0,k]})$, where $P(X^{(\nu)}_{j,[0,k]})$ is obtained via marginalization,
@@
where the index $\mu$ sums over $M_{2}$ trajectories sampled from $P_{0}$. What is the best choice for $P_{0}$? The ideal choice would be the conditional distribution $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]}^{(\nu)})$, as it makes the summand in Eq. 8 equal for all $\mu$, such that the variance of the estimate of $P(X_{j,[0,k]}^{(\nu)})$ is zero. However, this conditional distribution is not known a priori. We therefore generate $X_{i}$ and $X_{l}$ trajectories in the frozen field of $X_{j,[0,N]}^{(\nu)}$ (Fig. 1b) resulting in a distribution $P_{0}$ that is known analytically and is expected to be close to this conditional distribution. To exactly compensate for the remaining deviations of $P_{0}$ from the desired conditional distribution, we employ, in the spirit of Rosenbluth-Rosenbluth(RR)-PWS18, a reweighing of the $X_{i}^{(\mu)}$ and $X_{l}^{(\mu)}$ trajectories on-the-fly with weights proportional to the ratio of the two distributions. We resample the trajectories with these weights after every $\delta t$ time, meaning that we sample $M_{2}$ trajectories with replacement from the weighted trajectory ensemble (Fig. 1c). This procedure exactly generates $X_{i}^{(\mu)}$ and $X_{l}^{(\mu)}$ trajectories according to the desired conditional distribution $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]}^{(\nu)})$ (see EM and Supplemental Material (SM) section SM-C).
@@
To summarize, for each of the $M_{1}$ pairs of $X_{i,[0,N]}$ and $X_{j,[0,N]}$ trajectories, we simulate a joint ensemble of $M_{2}$ new $X_{i}$ and $X_{l}$ trajectories to estimate the denominator in the logarithm in Eq. 6, and a separate ensemble of $M_{2}$ new $X_{l}$ trajectories to estimate the numerator (Fig. 1). The trajectories in each ensemble are resampled on-the-fly with the RR scheme, giving both the numerator and denominator in Eq. 6 as Monte-Carlo averages. The computational cost thus scales as $2M_{1}M_{2}$. The transfer entropy estimate is unbiased18 and the statistical accuracy can be arbitrarily improved by increasing $M_{1}$ and $M_{2}$. A pseudocode for the algorithm is available in SM-B. Aside from Schreiber’s transfer entropy1, other trajectory-based metrics of directional information transfer such as directed information22, conditional transfer entropy23 and filtered transfer entropy24, also can be derived from conditional distributions of trajectories. Hence TE-PWS can be used to compute all such metrics at similar cost, as shown in SM-G.
@@
We demonstrate the validity of the method by analyzing two examples of an Ornstein-Uhlenbeck (OU) process with feedback for which transfer entropy rates are exactly available since the trajectories are Gaussian-distributed. Eq. 1 represents an OU process when it is linear with $\mathbf{F}=-\mathbf{a}\mathbf{X}$, where $\mathbf{a}$ is a spring constant matrix. The first example is a two-variable OU process (model A) for which the transfer entropy rates are analytically available (see EM). Additionally, since the trajectories are Gaussian distributed, each term in Eq. 3 can be exactly computed by estimating the covariance of trajectories, a commonly used approach called the Gaussian framework for the transfer entropy13, 14 (see SM-H). On comparing these values with TE-PWS in Fig. 2a, we find that TE-PWS gives accurate and unbiased estimates of $\dot{\mathcal{T}}_{X_{1}\to X_{2}}^{[k]}\equiv\mathcal{T}_{X_{1}\to X_{2}}^{[k]}/\delta t$ for every $k$, converging to $\dot{\mathcal{T}}_{X_{1}\to X_{2}}$ for large $k$. A second benchmark is a three-variable model (model B) for describing the stochastic dynamics of gene expression and the growth rate of bacteria25. The model is non-bipartite, meaning that the diffusion constant matrix is non-diagonal and the mutual information rates are not finite. Yet transfer entropy rates are finite and semi-analytically available24. As shown in Fig. 2b and c, the steady-state transfer entropy rate from TE-PWS converges to the known values. Details about the models are provided in SM-A.
@@
Figure 2: Convergence of transfer entropy rate estimates in linear models. (a) shows $\dot{\mathcal{T}}_{X_{1}\to X_{2}}^{[k]}\equiv\mathcal{T}_{X_{1}\to X_{2}}^{[k]}/\delta t$ in the two-variable OU process (model A) with feedback as a function of the history length $k+1$. The black line is from TE-PWS, the solid red line is the exact result from the Gaussian framework, and the red dashed-dot and dashed lines are analytical results for $\dot{\mathcal{T}}_{X_{1}\to X_{2}}^{[k=0]}$ (single-step) and $\dot{\mathcal{T}}_{X_{1}\to X_{2}}$, respectively. (b) and (c) show transfer entropy rates in the linear stochastic gene expression model (model B)25. Symbols are from TE-PWS and dashed lines are semi-analytical results.
Data processing inequality. We demonstrate the utility of TE-PWS by applying it to a three-node motif to study whether transfer entropies obey the Data Processing Inequality (DPI). Unidirectional flow of information between different nodes in a network leads to a DPI for the mutual information26. For a general three-variable process, if the flow of information is $X_{1}\to X_{2}\to X_{3}$, i.e., without feedback, $I(X_{1,[0,N]};X_{3,[0,N]}|X_{2,[0,N]})=0$; here the right arrows denote flow of information as mediated either via activation or repression. This leads by the chain rule to $I(X_{1,[0,N]};X_{3,[0,N]})\leq I(X_{1,[0,N]};X_{2,[0,N]})$26. As transfer entropy equals mutual information in the absence of feedback, it also obeys $\mathcal{T}_{X_{1}\to X_{3}}\leq\mathcal{T}_{X_{1}\to X_{2}}$. This relation bounds the amount of information that can be transmitted from input to output through an intermediate variable, yet is only valid in the absence of feedback. In the presence of an $X_{2}\rightarrow X_{1}$ feedback, i.e., $X_{1}\rightleftarrows X_{2}\to X_{3}$, the mutual information continues to obey its DPI since $X_{1}$ and $X_{3}$ remain independent conditional on $X_{2}$ (see SM-I), while the transfer entropy formally doesn’t27, 28. This can be rationalized by considering the limit $X_{1}\leftarrow X_{2}\rightarrow X_{3}$, where $X_{2}$ controls both $X_{1}$ and $X_{3}$. Here the $X_{1}$ trajectories would still be predictive of fluctuations in $X_{3}$, thus $\mathcal{T}_{X_{1}\to X_{3}}\geq 0$ even as $\mathcal{T}_{X_{1}\to X_{2}}=0$. However, where the crossover from the feedforward to the feedback-dominated regime occurs, and to what extent transfer entropies can overcome the DPI in practice, is currently not understood.
Figure 3: (a) and (b) are typical $X_{1}$ (green) and $X_{2}$ (black) trajectories for models C and D at $a_{12}=2a_{21}=0.4$ and $a_{12}=a_{21}=-6$, respectively. (c) Transfer entropy rates in model C, denoted as $\dot{\mathcal{T}}_{X_{1}\to X_{3}}^{(\mathrm{C})}$ and $\dot{\mathcal{T}}_{X_{1}\to X_{2}}^{(\mathrm{C})}$, as a function of an increasing ratio of feedback to feedforward strength $f^{*}=a_{12}/a_{21}$, where the feedforward strength is kept constant at $a_{21}=0.2$. (d) Similar to (c) but for model D, where the feedforward strength is kept constant at $a_{21}=-4$. (e) Ratio of transfer entropy rates overcomes DPI bound (dashed line at $\mathcal{T}^{*}=1$) with increasing feedback.
We implement the motif $X_{1}\rightleftarrows X_{2}\to X_{3}$ in two diffusive models of mutual repression between $X_{1}$ and $X_{2}$, labeled as models C and D. In both models, $X_{3}$ rapidly copies the state of $X_{2}$ such that the information loss from $X_{2}$ to $X_{3}$ is low. Model C is a three-dimensional OU process with linear feedback, where the ratio of the feedback to feedforward spring constants $f^{*}\equiv a_{12}/a_{21}$ is varied to study the violation of DPI. Model D is a nonlinear extension of C inspired by a genetic toggle switch29 where the drifts for $X_{1}$ and $X_{2}$ are changed to $F_{1}=-a_{11}X_{1}-a_{12}(1+X_{1}^{2})/(1+X_{1}^{2}+X_{2}^{2})$ and $F_{2}=-a_{22}X_{2}-a_{21}(1+X_{2}^{2})/(1+X_{1}^{2}+X_{2}^{2})$; we choose $a_{12}/a_{21}=f^{*}$ such that the two models can be compared. Other model parameters are provided in SM-A. Typical $X_{1}$ and $X_{2}$ trajectories at high feedback are shown in Figs. 3a and b. Model C merely shows regression to the mean for both $X_{1}$ and $X_{2}$, while model D additionally shows switching between (low,high) and (high,low) values of $(X_{1},X_{2})$ in the regime of both strong feedforward and feedback coupling, around $f^{*}=1$.
We plot in Figs. 3c and d transfer entropy rates for the two models. Each data point is obtained from steady state trajectories using TE-PWS. In model C, increasing the feedback strength $f^{*}$ keeps $\dot{\mathcal{T}}_{X_{1}\to X_{2}}$ unchanged while $\dot{\mathcal{T}}_{X_{1}\to X_{3}}$ monotonically increases. The former follows analytically from the expression for the transfer entropy in a two-dimensional OU process24 (see EM). The latter is an empirical result that arises because, when $f^{*}$ is increased, $X_{1}$ becomes increasingly correlated with $X_{2}$ while $X_{2}$ continues to be copied accurately by $X_{3}$. In contrast, in model D, both $\dot{\mathcal{T}}_{X_{1}\to X_{2}}$ and $\dot{\mathcal{T}}_{X_{1}\to X_{3}}$ peak near the switching regime, phenomenologically similar to how a bistable three-node motif with feedback has been shown to behave through approximate theory15. The values of the transfer entropy rates in model D are also amplified many-fold compared to model C due to stronger correlations between $X_{1}$ and $X_{2}$ resulting from the stronger, nonlinear coupling, with the concomitant switching.
In Fig. 3e we have plotted the degree of DPI violation in both models, quantified as the ratio of transfer entropy rates $\mathcal{T}^{*}\equiv\dot{\mathcal{T}}_{X_{1}\to X_{3}}/\dot{\mathcal{T}}_{X_{1}\to X_{2}}$, as a function of the ratio of the feedback to feedforward strengths, $f^{*}$. Surprisingly, we find that regardless of the nature of the variation of the individual transfer entropy rates with increasing feedback, i.e., monotonic or non-monotonic (Figs. 3c and d), the ratio $\mathcal{T}^{*}$ monotonically increases with increasing feedback (Fig. 3 e). Moreover, the ratio overcomes the DPI bound when the strength of the feedback becomes comparable to that of the feedforward coupling. Our results thus show that when the feedback $X_{2}\to X_{1}$ dominates over the feedforward interaction, $X_{1}\to X_{2}$, the feedforward entropy $\mathcal{T}_{X_{1}\to X_{3}}$ becomes larger than the feedforward entropy $\mathcal{T}_{X_{1}\to X_{2}}$. The mutual information, on the other hand, continues to obey its DPI, $I(X_{1,[0,N]};X_{3,[0,N]})\leq I(X_{1,[0,N]};X_{2,[0,N]})$ (see SM-I). We expect further analytical work in the OU process to be able to support this empirical result. Moreover, as the transfer entropy has been shown to limit the costs and benefits of information transmission6, 7, we expect the violation of the DPI to have significant physical consequences in functional networks.
Accuracy of approximate methods. An exact method is essential for the rigorous testing of the accuracy of approximate methods. This is now possible for the first time with our technique. We have compared the ground-truth TE-PWS results with those obtained from the Gaussian framework and the KSG algorithm30, in the linear model A and in the nonlinear model D (see Table 1 in EM). We find that all approximate methods incur large systematic errors depending on the model. The Gaussian framework as expected gives accurate results in the linear model but fails in the nonlinear model. $\dot{\mathcal{T}}^{[k]}$ from the KSG method in the limit $(k\to\infty)$ has a systematic error due to a) its assumption of local uniformity of the distribution of the data points and b) the downsampling that becomes necessary to obtain a converged estimate (see SM-H). To avoid the systematic error, KSG is often truncated at only the $k=0$ term of Eq. 3, yielding the so-called single-step transfer entropy rate1, 31, 11, 24. As the table shows and is demonstrated in Fig. 2a, the single-step transfer entropy rate is in general not an accurate estimate of the long-time limit. Interestingly, as the table shows, TE-PWS is not only exact but also highly cost efficient, being either comparable or orders of magnitude cheaper than approximate methods. This is because a) TE-PWS can calculate the transfer entropy for all $k$ in one run, and hence can average over all $k$ in the long time limit as shown in Eq. 2, and b) for any given $k$, the cost for TE-PWS scales proportional to $k$, while for Gaussian and KSG the scaling is, respectively, $\sim k^{3}$ (see SM-H) and superlinear due to the cost of nearest neighbor search32, 33 .
In conclusion, we have developed a method that for the first time makes it possible to compute transfer entropies exactly for any stochastic model. We expect transfer entropies computed by TE-PWS to be used as ground truth for a wide range of goals, such as the characterization and design of information flow in natural and engineered information processing systems, and causality detection.
Data availability Python implementations of TE-PWS for diffusive and jump processes, and data that reproduce the findings of this study, are openly available on Zenodo at https://zenodo.org/doi/10.5281/zenodo.1361736534.
Appendix A End Matter
where $r_{2}=[a_{11}^{2}+(D_{11}/D_{22})a_{21}^{2}-2(D_{12}/D_{22})a_{11}a_{21}]^{1/2}$. On the other hand, if Eq. 3 is truncated at the first term after starting from a steady-state, the resultant approximation, called the single-step transfer entropy rate, is given by24
$\lim_{\delta t\to 0}\frac{1}{\delta t}\mathcal{T}_{X_{1}\to X_{2}}^{[0]}=\frac{1}{4D_{22}}\frac{a_{21}^{2}|\bf{V}|}{V_{22}}$ (11)
where $\bf{V}$ is the stationary covariance matrix with elements $V_{ij}\equiv\langle X_{i}(0)X_{j}(0)\rangle$, analytical expressions for which can be found in [24]. In general, depending on parameter values such as feedback and noise strengths, Eqs. 10 and 11 yield different values. Specifically, Eqs. 10 and 11 have been used to plot the two distinct dashed lines in Fig. 2a. Additionally, we note that for model C in the section Data processing inequality of the main text, the dynamics of $X_{1}$ and $X_{2}$ comprises that of a two-dimensional OU process because there is no feedback from $X_{3}$, only between $X_{1}$ and $X_{2}$. Hence the transfer entropy rate $\dot{\mathcal{T}}_{X_{1}\to X_{2}}$ is given by Eq. 10, which does not depend on $X_{2}\to X_{1}$ feedback strength $a_{12}$, i.e., on $f^{*}$, as stated in the main text (see Fig. 3c).